OpenStudy (anonymous):

Evaluate the limit. As h-->0, the square root of a+h - square root of a all over h.

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \sqrt{a+h}-\sqrt{a}/h\]

OpenStudy (anonymous):

 evaluate the limit : \[(1/2\sqrt{a})\]